A point charge of +7.5µ C is located 0.31 m from a second point charge of -6.8µ C. What is the absolute magnitude of the force on each charge?

(please show work! thanks!)
Answer #1

Force on each charge =

F = 9*109 * 7.5*10-6*6.8*10-6/(0.312)

= 4.776N
Answer #2
F = kq1q2/r^2 = 4.7719N
Answer #3
F = k(q10*(q2)/r^2

F = (7.5*6.8)*(9*10^9)*(10^-12)/(0.31)^2 = 4.77 N